OpenStudy (anonymous):

how do I do this question 2/x-3 >_2 please help me on thisthanks :D

OpenStudy (anonymous):

is that (2/x) - 3, or 2/(x-3)? and is that -2?

OpenStudy (anonymous):

no is an greater equal sign

OpenStudy (anonymous):

and my other question?

OpenStudy (anonymous):

nop is 2 divided by x-3

OpenStudy (anonymous):

\[\frac{2}{x-3} \ge 2\] \[2 \ge 2x - 6\] \[8 \ge 2x\] \[4 \ge x\]

OpenStudy (anonymous):

hey there how did u do that ?

OpenStudy (anonymous):

multiplied both sides by x-3 and solved. Notice the solution is only valid for \[3 \lt x \le 4\] because at x=3 the equation is undefined, and at x < 3, the solution is negative.

OpenStudy (anonymous):

the graph: http://bit.ly/pauLUV

OpenStudy (anonymous):

omq thanks so much :)

OpenStudy (anonymous):

hey the solution I have to give in interval notation so how do u think I should do it?

OpenStudy (anonymous):

(3,4]

OpenStudy (anonymous):

yeah Thats what I got :)

OpenStudy (anonymous):

thankss

OpenStudy (anonymous):

hey did you know how to solve (5-3/x)/(1+4/x^2)